Find the Difference Quotient
f(x)=x^2 -2x +9

Answer by fractalier(6550)

For f(x)= x^2 - 2x + 9
[f(x + h) - f(x)] / h =
[((x+h)^2 - 2(x+h) + 9) - (x^2 - 2x + 9)] / h =
[(x^2 + 2hx + h^2 - 2x - 2h + 9) - (x^2 - 2x + 9) / h =
(2hx + h^2 - 2x) / h =
2x + h - 2 =
2x - 2 + h